# SIRD: A Sender-Informed, Receiver-Driven Datacenter Transport Protocol

### v4

@@ 2.2. The impact of ASIC trends on buffering
Congestion control protocols generally depend on buffering to offset coordination and control loop delays and, as a result, face a throughput-buffering trade-off. Maximum bandwidth utilization can trivially be achieved with high levels of in-network buffering, but at the cost of queuing-induced latency and expensive dropped packet retransmissions. Conversely, low buffering can lead to throughput loss for protocols that are slow in capturing newly available bandwidth.
@@
High-speed packet buffering is handled by low latency and relatively small SRAM buffers in switch ASICs. Figure 1 shows that the size of these buffers is not increasing in size as fast as the bisection bandwidth of switch ASICs. For example, nVidia’s top-end Spectrum 4 ASIC (rightmost point in Figure 1) has a $160$MB buffer, which corresponds to $3.13$MB per Tbps of bisection bandwidth (Nvidia, 2022). The previous $12.8$Tbps and $6.4$Tbps top-end Spectrum ASICs were equipped with $5$MB and $6.6$MB per Tbps respectively (Mellanox, 2020, 2017). Unfortunately, future scaling of SRAM densities appears unlikely given CMOS process limitations (Chang et al., 2022; David Schor, 2022). In parallel, datacenter round-trip times (RTT) are not falling as they are dominated by host software processing, PCIe latency, and ASIC serialization latencies. Consequently, CC protocols must handle higher BDPs with less switch buffer space at their disposal.
@@
To better absorb instantaneous bursts of traffic, switch packet buffers are generally shared among egress ports. Some ASICs advertise fully-shareable buffers while others implement separate pools or statically apportion some of the space to each port (Ghabashneh et al., 2022). On top of the physical implementation, various buffer-sharing algorithms dynamically limit each port’s maximum allocation to avoid unfairness (Alizadeh et al., 2010). However, if a large part of the overall buffer is occupied, the per-port cap becomes more equitable, limiting burst absorbability (Ghabashneh et al., 2022).
Figure 2 provides some insight into whether existing congestion control schemes are future-proof given these hardware trends by measuring Homa’s buffer occupancy distribution as a function of load for the Websearch workload (Alizadeh et al., 2013). Hosts generate flows following the Poisson distribution but traffic patterns are often more bursty (Zhang et al., 2017) and thus challenging (full methodology in §5.1). To achieve best-in-class message latency and link utilization Homa uses Ethernet priorities and overcommitment of downlinks, the latter drastically increasing queuing in the fabric. Figure 2 compares the buffer occupancy distribution (assuming infinite buffers in the experiment) with the actual resources of two leading ASICs, adjusted to the port radix of the simulation. The simulated ToR switches have $16$ $100$Gbps downlinks to servers (Alizadeh et al., 2013) and $4$ $400Gbps$ uplinks for a total bisection bandwidth of $3.2$Tbps. Using the Spectrum 4 buffer-bandwidth ratio of $3.13$MB per Tbps corresponds to $0.31$MB of buffer per $100Gbps$ port (Figure 2 left) and a $10$MB total ToR buffer (right). In practice, the buffer is neither partitioned nor fully shared; however, between the two extremes, Homa operates with more than the per-port buffer and is close to overflowing a fully shared buffer. In the latter case, the presence of persistent queuing also limits the flexibility of the buffer-sharing algorithm.
@@
Figure 2. Queuing CDFs across downlinks and ToRs for Homa under increasing network load under the Websearch workload (Alizadeh et al., 2013). The dotted lines represent the switch buffer size adjusted to the radix of our simulated ToR.
3. SIRD Design Pillars
@@
SIRD is an end-to-end receiver-driven scheme that manages exclusive links proactively and shared links reactively. In practice, this means that receivers perform precise credit-based scheduling when the bottleneck is their downlinks and use well-established congestion signals and algorithms to coordinate over shared links. SIRD does not face the limitations discussed in §2 as it:
@@
  • causes minimal buffering in the network by largely eliminating the need to overcommit downlinks.
  • does not need Ethernet priorities to deliver messages with low latency, thanks to minimal buffering.
@@
Efficient credit allocation: SIRD’s objective is to allow receivers to send the appropriate amount of credit depending on the real time availability of the bottleneck link. For downlinks the task is trivial as all the link’s capacity is managed by one receiver. However, when the bottleneck is shared (sender uplink or core link), SIRD receivers detect competition for bandwidth and adjust the amount of issued credit dynamically. For example, if a sender is the bottleneck because it transmits continuously to two receivers, each of them should converge to allocating half a BDP worth of credit. This approach makes the distribution of credit efficient because it is allocated to senders that can promptly use it rather than being accumulated by congested senders.
@@
SIRD implements efficient credit allocation through informed overcommitment. Each receiver is allotted a limited amount of available credit $B$, and the objective is to distribute it to senders that can use it. The minimum valid value of $B$ is $1\times BDP$ as this is the amount of credit required to pull an equal amount of inbound traffic and fully utilize the downlink. Higher values of $B$ lead to downlink overcommitment. Unlike Homa’s controlled overcommitment that is statically configured and credits the $k$ senders sending the shortest messages, SIRD aggressively reduces overcommitment because it redirects credit to senders that can use it, based on feedback from senders and the network.
@@
Figure 3. Average buffering at ToRs versus maximum achieved goodput when sweeping the overcommitment parameter for SIRD (informed overcommitment) and Homa (controlled overcommitment). The former is a continuous parameter while the latter is discrete. Results obtained in simulation by injecting application traffic at 95Gbps/link across 144 100Gbps servers based on the Websearch (Alizadeh et al., 2013) workload (WKc); see §5.1 for the full methodology.
@@
Figure 3 showcases the benefit of efficient credit allocation in SIRD by comparing the throughput-buffering trade-off across equivalent overcommitment levels for Homa ($k$) and SIRD ($B$) under high network load, when CC is essential. SIRD’s informed overcommitment matches Homa’s goodput with 14×14$\times$14 × less overcommitment and 13×13$\times$13 × less queuing.
@@
SIRD implements informed overcommitment through a control loop at each receiver, which dynamically adjusts credit allocation across senders based on sender and network feedback. When a sender is concurrently credited by multiple receivers, the sender receives credit faster than it can consume it, causing it to accumulate. This is undesirable as receivers have a limited amount of credit to distribute. Equivalently, limiting accumulation makes credit available to the senders that can use it. SIRD’s control loop achieves this by setting an accumulated credit threshold, $SThr$, that is similar in spirit to DCTCP’s marking threshold (Alizadeh et al., 2010), Swift’s target delay(Kumar et al., 2020), or HPCC’s $\eta$ (Li et al., 2019). At each sender, when the amount of accumulated credit across all receivers exceeds $SThr$, a bit is set in all outgoing data packets. Based on arriving bit values, each receiver executes DCTCP’s well established AIMD algorithm and reduces the amount of credit allocated to the congested sender. SIRD can also use other algorithms or signals such as end-to-end delay on infrastructures with precise timestamping support (Kumar et al., 2020), In-Band Telemetry (Li et al., 2019; Wang et al., 2023) if available.
@@
SIRD applies the same approach to handle core congestion using ECN as its signal, similar to DCTCP. Receivers observe whether the CE bit of data packets is set and execute the DCTCP algorithm to adjust per-sender credit limits. Similar to Swift (Kumar et al., 2020), each receiver runs two AIMD algorithms in parallel, one for senders and one for the core, and then enforces the decision of the most congested one. This separation allows the differentiation of the signal or the algorithm given the particular constraints or capabilities of hosts and network switches. In this paper, we use the same proven and simple combination for both and leave the exploration of more complex algorithms and signals to future work.
@@
Starting at line rate: Prior work (Montazeri et al., 2018; Gao et al., 2015; Hu et al., 2020; Handley et al., 2017) has demonstrated the latency benefits of starting transmission at line rate, without a gradual ramp-up or preceding control handshakes that take at least one RTT. SIRD’s credit allocation mechanism operates while transmission progress is made, and thus, senders immediately start at line rate, sending the first BDP bytes unscheduled (without waiting for credit).
@@
We further optimize the design based on the following simple observation: small messages benefit the most from unscheduled transmission since their latency is primarily determined by the RTT while throughput-dominated messages see minimal gain. For example, in the absence of queuing, delaying transmission by one RTT increases the end-to-end latency of a message sized at $10\times BDP$ by 9% compared to 200% for a single-packet message. Therefore, to reduce unnecessary bursty traffic and queuing, SIRD senders do not start transmitting messages larger than a configurable threshold (UnschT) before explicitly receiving credits from the receiver. For messages smaller than UnschT, senders send the first $min(BDP,msg\_size)$ bytes without waiting for credit. For example, if $UnschT=2\times BDP$ and $msg\_size=1.5\times BDP$, the first $BDP$ is unscheduled and the other $0.5$ is scheduled.
Ethernet priorities are not required: Homa’s main contribution is to leverage Ethernet priorities to solve for latency in a system that aggressively overcommits without sender state information. By efficiently allocating credit, SIRD maximizes utilization with minimal buffering, thereby largely eliminating its impact on latency, and removing the requirement for priorities. If available, SIRD can optionally use one separate priority lane to slightly lower the RTT of control packets and to deliver small messages even faster (see §5.5 for analysis on priority sensitivity).
4. SIRD Design
@@
At a high level, SIRD is an RPC-oriented protocol, similar to other recently proposed datacenter transports (Montazeri et al., 2018; Gao et al., 2015; Cai et al., 2022). SIRD may be used to implement one-way messages or remote procedure calls (Montazeri et al., 2018; Kalia et al., 2019; Kogias et al., 2019). SIRD assumes that the length of each message is known or that data streams will be chunked into messages. SIRD further assumes that ECN is configured in all network switches; the ECN threshold is to be set according to DCTCP best practices (Alizadeh et al., 2010). SIRD is designed to be layered on top of UDP/IP for compatibility with all network deployments. The UDP source port is randomly selected for each packet for fine-grain load balancing, allowing an ECMP network to behave as efficaciously as random packet spraying (Dixit et al., 2013). We make no assumption of lossless delivery, network priorities, or smart NICs and switches.
In the rest of this section we present SIRD focusing on the mechanisms that enable informed overcommitment.
4.1. Base Sender-Receiver Interaction
SIRD defines three main packet types:
  1. (1)
    DATA: a packet that contains part of a message payload. There are two types of DATA packets: scheduled, that are sent after receiving credit, and unscheduled.
  2. (2)
    CREDITREQ: a control packet sent by a sender to a receiver to indicate that the former wants to send scheduled data, either an entire message if $msg\_size>UnschT$, or else just the scheduled portion of a message smaller than $UnschT$.
  3. (3)
    CREDIT: a control packet sent by the receiver to the sender to schedule the transmission of one DATA packet.
@@ 4.1. Base Sender-Receiver Interaction
Generally, the flow of packets consists of credit flowing from receivers to senders and data flowing in the opposite direction. Credit leaves the receiver in a CREDIT packet, is used by the sender, and returns in a scheduled DATA packet.
4.2. Credit Management
@@
The global credit bucket $B$ controls overcommitment by capping the total number of outstanding credits that a receiver can issue. Configuring $B\geq BDP$ is necessary to ensure maximum throughput. Further, $B$ bounds the queuing length from scheduled messages to $B-BDP$ bytes at the ToR’s downlink. Our experiments show that a value of $B$ as low as $1.5\times BDP$ is sufficient to ensure high link utilization in challenging workloads.
@@
Unlike protocols that manage congestion on a per-flow basis, SIRD does so per sender. In SIRD, all flows will face the same level of congestion since it performs packet spraying. Thus, SIRD receivers aggregate congestion signals from all flows and can react faster. From an implementation perspective, receivers need to keep less state, which is beneficial especially for a hardware implementation.
4.3. Informed Overcommitment
@@
The informed overcommitment control loop uses two input signals that communicate the extent of congestion in the network and at senders. The signals are carried in DATA packets and are the ECN bit in the IP header set by the network and a bit in the SIRD header set by the sender. Each receiver runs two separate AIMD (additive-increase, multiplicative decrease) control loops and uses the most conservative of the two (similar to Swift (Kumar et al., 2020)) to adjust per-sender credit bucket sizes. Each control loop is configured with its own marking threshold ($NThr$ and $SThr$). $NThr$ should be set according to DCTCP best practices (Alizadeh et al., 2010) to limit queuing in the network core. Note that $NThr$ is much higher than the allowed persistent queueing at ToR switches as controlled by $B$. Thus, ToR switches never have to mark ECN.
@@
$SThr$ should be set to limit the amount of credit a sender can accumulate, thus allowing receivers to efficiently distribute their limited aggregate credit. Intuitively, $SThr$ determines the level of accumulated credit the control loop is targeting when a sender is congested, i.e., receiving credit faster than it can use it. Setting it too high means each sender can accumulate substantial amounts of credit, and consequently, B would need to be configured higher to increase aggregate credit availability. Conversely, a very low value of SThr does not allow the control loop any slack when converging to a stable state and can cause throughput loss.
@@
To understand the relationship between $SThr$ and $B$, we analytically examine the congested-sender scenario from the perspective of a receiver R to find: how much total credit (B) R needs when receiving from $k$ congested senders to still have enough available to be able to saturate its downlink. Assuming each congested sender sends to $f$ receivers in total, the available uplink bandwidth for R is $BW/f$. Assuming uniform link speeds, we are interested in the case where senders are the bottleneck and cannot saturate R, or:
(1) $BW_{supply\_k}<BW_{demand\_R}=>\frac{BW}{f}<\frac{BW}{k}=>f>k$
@@
In this case, each of the $k$ senders accumulates up to $SThr$ credit in stable state, or $SThr/f$ from each receiver assuming an equal split (see §4.5). Therefore, R’s B must be large enough to allow $1xBDP$ of credit to be in flight (for any new senders) despite accumulation at congested senders, or:
(2) $B\geq BDP+\sum_{k}{\frac{SThr}{f}\;;\;f\geq 2,f>k}$
@@
If $k_{max}$ is the number of congested senders that maximizes the sum term, then $f_{max}=k_{max}+1$ (denominator) and the maximum value of the term is:
(3) $\sum_{k}^{k_{max}}{\frac{SThr}{k_{max}+1}}=\frac{SThr}{k_{max}+1}\sum_{k}^{k_{max}}{1}=SThr\frac{k_{max}}{k_{max}+1}<SThr$
@@
It follows that, in steady state, R can account for any number of congested senders as long as $B\geq BDP+SThr$. Under dynamic traffic patterns (see §5.5) higher values of B can help increase overall utilization. Further, policies other than fair sharing can loosen this property. For example, if senders prioritize some receivers, then, in the worst case where R is de-prioritized by all k senders, Equation 2 loses $f$ from the denominator and B depends on the number of worst-case congested senders.
Table 1. Core configuration parameters.
@@
2:procedure onDataPacket($pkt,i$$\triangleright$ i:sender
3:    $credit\leftarrow{getCredit}(pkt)$
4:    $b\leftarrow b-credit$
5:    $sb_{i}\leftarrow sb_{i}-credit$
6:    $senderBkt_{i}\leftarrow{Sender_{AIMD}}(senderBkt_{i},pkt.sird.csn)$
7:    $netBkt_{i}\leftarrow{Net_{AIMD}}(netBkt_{i},pkt.ip.ecn)$
8:end procedure
9:procedure sendCredit() $\triangleright$ Runs when $b+min(rem_{i},MSS)\leq B$
10:     $senderList\leftarrow activeSenders.filter($ $sb_{i}+min(rem_{i},MSS)\leq min(senderBkt_{i},netBkt_{i}))$
11:    $s\leftarrow{policySelect}(senderList)$
12:    $credit\leftarrow min(rem_{i},MSS)$
13:    ${sendCredit}(s,credit)$
14:    $b\leftarrow b+credit$; $sb_{i}\leftarrow sb_{i}+credit$; $rem_{i}\leftarrow rem_{i}-credit$
15:end procedure

### v5

@@ 2.2 The impact of ASIC trends on buffering
Congestion control protocols depend on buffering to offset coordination and control loop delays and, as a result, face a throughput-buffering trade-off. Maximum bandwidth utilization can trivially be achieved with high levels of in-network buffering, but at the cost of queuing-induced latency and expensive dropped packet retransmissions. Conversely, low buffering can lead to throughput loss for protocols that are slow in capturing newly available bandwidth.
@@
High-speed packet buffering is handled by small SRAM buffers in switch ASICs, the size of which is not increasing as fast as bisection bandwidth. For example, nVidia’s top-end Spectrum 4 ASIC has a $160$MB buffer, which corresponds to $3.13$MB per Tbps of bisection bandwidth [58]. The previous $12.8$Tbps and $6.4$Tbps top-end Spectrum ASICs were equipped with $5$MB and $6.6$MB per Tbps respectively [53, 52] (full list in Appendix A). Unfortunately, future scaling of SRAM densities appears unlikely given CMOS process limitations [17, 24]. In parallel, datacenter round-trip times (RTT) are not falling as they are dominated by host software processing, PCIe latency, and ASIC serialization latencies. Consequently, as link speeds increase, CC protocols must handle higher BDPs with less switch buffer space at their disposal.
@@
To better absorb instantaneous bursts of traffic, switch packet buffers are generally shared among egress ports. Some ASICs advertise fully shareable buffers while others implement separate pools or statically apportion some of the space to each port [33]. On top of the physical implementation, various buffer-sharing algorithms dynamically limit each port’s maximum allocation to avoid unfairness [8]. However, if a large part of the overall buffer is occupied, the per-port cap becomes more equitable, limiting burst absorbability [33].
Figure 1 provides some context by comparing the buffer use of Homa, a state-of-the art RD protocol, with recent switch buffer capacities under the Websearch workload [10]. The dotted lines represent the per-port (left) and completely shared (right) buffer capacities of the hypothetical switches, adjusted on a per-unit basis to the port radix of our simulation methodology (§6.2). In practice, the buffer is neither partitioned nor fully shared; however, between the two extremes, hardware trends are constraining the throughput-buffering trade-off.
@@
Figure 1: Homa queuing CDFs under various network loads for workload Websearch [10]. The dotted lines represent the switch buffer size adjusted to the actual radix of our simulated ToR; see §6.2.
3 SIRD Design Pillars
@@
SIRD is an end-to-end receiver-driven scheme that manages exclusive links proactively and shared links reactively. In practice, this means that receivers perform precise credit-based scheduling when the bottleneck is their downlinks and use congestion feedback to coordinate over shared links. SIRD does not face the limitations of prior work as it:
@@
  • causes minimal buffering in the network by drastically reducing the need for downlink overcommitment.
  • does not need switch priority queues to deliver messages with low latency, thanks to minimal buffering.
@@
Efficient credit allocation: SIRD’s design allows receivers to send the appropriate amount of credit depending on the real-time availability of the bottleneck link. For downlinks the task is trivial as the link’s capacity is managed by one receiver. However, when the bottleneck is shared (sender uplink or core link), SIRD receivers detect competition for bandwidth and adjust the amount of issued credit dynamically. For example, if a sender is the bottleneck for two receivers, each will allocate the appropriate amount of credit for their share of the sender’s bandwidth. This approach makes the distribution of credit efficient because it is given to hosts that can promptly use it rather than being accumulated by congested senders. As a result, SIRD can achieve high link utilization using a small amount of credit, i.e., with little overcommitment.
@@
SIRD combines efficient credit allocation and overcommitment in informed overcommitment. Each receiver is allotted a limited amount of available credit $B$. The minimum valid value of $B$ is $1\times BDP$ as this is the amount of credit required to pull $1\times BDP$ of inbound traffic and fully utilize the downlink. Higher values of $B$ lead to downlink overcommitment, which increases tolerance to congested shared links but also increases expected queuing.
@@
Figure 2: Mean buffering at ToRs versus maximum achieved goodput when sweeping the overcommitment parameter for SIRD (informed overcommitment) and Homa (controlled overcommitment). Results obtained in simulation by running the Websearch workload at 95Gbps on 100Gbps links across 144 servers; see §6.2.
@@
Figure 2 highlights the benefit of efficient credit allocation in SIRD by comparing the throughput-buffering trade-off across equivalent overcommitment levels for Homa ($k$), which introduced and uses RD overcommitment, and SIRD ($B$) under maximum network load. Informed overcommitment yields the same overall goodput despite overcommitting downlinks 14×14$\times$14 × less, which leads to 13×13$\times$13 × lower queuing. The goodput-queuing trade-off benefit is maximum at high network load, under which congestion and loss are most likely, and naturally fades as load lessens.
@@
SIRD implements informed overcommitment through a control loop at each receiver, which dynamically adjusts credit allocation across senders based on sender and network feedback. When a sender is concurrently credited by multiple receivers, the sender receives credit faster than it can consume it, causing it to accumulate. This is undesirable as receivers have a limited amount of credit to distribute. Equivalently, limiting accumulation makes credit available to other senders that can actually use it. SIRD’s control loop rebalances credit by setting an accumulated credit threshold, $SThr$, that is similar in spirit to DCTCP’s marking threshold [8], Swift’s target delay[44], or HPCC’s $\eta$ [47]. At each sender, while the amount of accumulated credit from all receivers exceeds $SThr$, a bit is set in all outgoing data packets. Based on arriving bit values, each receiver adjusts the maximum amount of credit that can be allocated to the congested sender in an additive-increase/multiplicative-decrease (AIMD) manner. In this paper, we use DCTCP’s [8] AIMD algorithm.
SIRD uses the same mechanisms to handle core congestion with ECN as the signal. Receivers read the CE bit of data packets and adjust per-sender credit limits using AIMD.
@@
Each receiver runs two AIMD algorithms in parallel, one for senders and one for the core, each with its own input signal. The algorithm instance of the most congested area determines credit allocation. This separation of signals allows SIRD to use different signal/algorithm combinations to handle bottlenecks at hosts and switches. Beyond ECN, SIRD can use signals such as end-to-end delay on infrastructures with timestamping support [44] or In-Band Telemetry [47, 69]. In this paper, we use the same proven and simple combination for both and leave the exploration of more complex algorithms and signals to future work.
@@
Starting at line rate: Prior work [56, 31, 39, 36] has demonstrated the latency benefits of starting transmission at line rate, without a gradual ramp-up or preceding control handshakes that take at least one RTT. SIRD’s credit allocation mechanism operates while transmission progress is made, and thus, senders immediately start at line rate, sending the first BDP bytes unscheduled (without waiting for credit).
@@
We further optimize the design based on the following simple observation: small messages benefit the most from unscheduled transmission since their latency is primarily determined by the RTT while throughput-dominated messages see minimal gain. For example, in the absence of queuing, delaying transmission by one RTT increases the end-to-end latency of a message sized at $10\times BDP$ by 9% compared to 200% for a single-packet message. Therefore, to reduce unnecessary bursty traffic and queuing, SIRD senders do not start transmitting messages larger than a configurable threshold (UnschT) before explicitly receiving credits from the receiver. For messages smaller than UnschT, senders send the first $min(BDP,msg\_size)$ bytes without waiting for credit.
Switch priority queues not required: By efficiently allocating credit, SIRD achieves high link utilization with minimal fabric queuing, thereby largely eliminating its impact on latency. Thus, SIRD can be deployed without any in-network QoS support, though it can benefit in terms of tail latency from having a total of two priority levels.
4 SIRD Design
@@
At a high level, SIRD is an RPC-oriented protocol, similar to other recently proposed datacenter transports [56, 31, 16]. SIRD may be used to implement one-way messages or remote procedure calls [56, 42, 43]. SIRD assumes that the length of each message is known or that data streams will be chunked into messages. SIRD further assumes that ECN is configured in all network switches, with the ECN threshold set according to DCTCP best practices [8]. SIRD is designed to be layered on top of UDP/IP for compatibility with all network deployments. The UDP source port is randomly selected for each packet for fine-grain load balancing, allowing an ECMP network to behave as effectively as random packet spraying [26]. We make no assumption of lossless delivery, in-network priority queues, or smart NICs and switches.
SIRD defines two main packet types: (1) DATA: a packet that contains part of the payload of a message. Data packets can be scheduled, requiring credit to be sent, or unscheduled. Messages with size $>UnschT$ are scheduled and make an initial credit request by sending a zero-length DATA packet; (2) CREDIT: a control packet sent by the receiver to the sender to schedule the transmission of scheduled DATA.
@@
Generally, the flow of packets consists of credit flowing from receivers to senders and data flowing in the opposite direction. Credit leaves the receiver in a CREDIT packet, is used by the sender, and returns with a scheduled DATA packet.
4.1 Credit Management
@@
The global credit bucket $B$ controls overcommitment by capping the total number of outstanding credits that a receiver can issue. Configuring $B\geq BDP$ is necessary to enable link saturation. Further, $B$ bounds the queuing length from scheduled packets to $B-BDP$ bytes at the ToR’s downlink. Our analysis (§4.2) and evaluation (§6) show that setting $B$ as low as $1.5\times BDP$ is sufficient for high link utilization.
@@
4.2 Informed Overcommitment
@@
Informed overcommitment uses two input signals that communicate the extent of congestion in the network and at senders. The signals are carried in DATA packets and are the ECN bit in the IP header set by the network and a bit in the SIRD header set by the sender. Each receiver runs two separate AIMD (additive-increase, multiplicative decrease) control loops and uses the most conservative of the two (similar to Swift [44]) to adjust per-sender credit bucket sizes. Each control loop is configured with its own marking threshold ($NThr$ and $SThr$). $NThr$ should be set according to DCTCP best practices [8] to limit queuing in the network core. Note that $NThr$ is much higher than the maximum allowed $B-BDP$ of persistent queueing at ToR switches. Thus, ToR switches never have to mark ECN.
@@
$SThr$ should be set to limit the amount of credit a sender can accumulate, thus allowing receivers to efficiently distribute their limited aggregate credit. Intuitively, $SThr$ determines the level of accumulated credit the control loop is targeting when a sender is congested, i.e., receiving credit faster than it can use it. Setting it too high means each sender can accumulate substantial amounts of credit, and consequently, $B$ would need to be configured higher to increase aggregate credit availability. Conversely, setting $SThr$ too low does not allow the control loop any slack when converging to a stable state and can cause throughput loss.
@@
To understand the relationship between $SThr$ and $B$, we analytically examine a simple congested-sender scenario from the perspective of a receiver $R$ to find: how much total credit ($B$) does $R$ need when receiving from $k$ congested senders, to still have enough credit available to saturate its downlink. Assuming each congested sender sends to $f$ receivers in total, its available uplink bandwidth for $R$ is $BW/f$. Assuming uniform link speeds, we are interested in the case where senders are the bottleneck and cannot saturate $R$, or:
$BW_{supply\_k}<BW_{demand\_R}=>\frac{BW}{f}<\frac{BW}{k}=>f>k$ (1)
@@
In this case, each of the $k$ senders accumulates up to $SThr$ credit in stable state, or $SThr/f$ from each receiver assuming an equal split (see §4.4). Therefore, to be able to saturate its downlink, $R$’s $B$ must be large enough to allow $1xBDP$ of credit to be in flight despite accumulation at congested senders, or:
$B\geq BDP+\sum_{k}{\frac{SThr}{f}\;;\;f\geq 2,f>k}$ (2)
@@
If $k_{max}$ is the number of congested senders that maximizes the right side of the inequality, then $f_{max}=k_{max}+1$ ($f$ in denominator) and the maximum value of the term is:
$\sum_{k}^{k_{max}}{\frac{SThr}{k_{max}+1}}=\frac{SThr}{k_{max}+1}\sum_{k}^{k_{max}}{1}=SThr\frac{k_{max}}{k_{max}+1}<SThr$ (3)
@@
It follows that, in steady state, $R$ can account for any number of congested senders as long as $B\geq BDP+SThr$. Under dynamic traffic patterns (see §6.2.4) higher values of B can help increase the supply of credit in transient phases. Further, policies other than fair sharing can loosen this property. For example, if senders strictly prioritize some receivers, then, in the worst case where $R$ is de-prioritized by all k senders, Equation 2 loses $f$ from the denominator and B depends on the number of worst-case congested senders.
Table 1: Core configuration parameters.
@@
1:procedure onDataPacket($pkt,i$$\triangleright$ i:sender
2:    $credit\leftarrow{getCredit}(pkt)$
3:    $b\leftarrow b-credit$
4:    $sb_{i}\leftarrow sb_{i}-credit$
5:    $senderBkt_{i}\leftarrow{Sender_{AIMD}}(senderBkt_{i},pkt.sird.csn)$
6:    $netBkt_{i}\leftarrow{Net_{AIMD}}(netBkt_{i},pkt.ip.ecn)$
7:end procedure
8:procedure onSendCreditTick() $\triangleright$ Runs when $b+min(rem_{i},MSS)\leq B$
9:     $senderList\leftarrow activeSenders.filter($ $sb_{i}+min(rem_{i},MSS)\leq min(senderBkt_{i},netBkt_{i}))$
10:    $s\leftarrow{policySelect}(senderList)$
11:    $credit\leftarrow min(rem_{i},MSS)$
12:    ${sendCredit}(s,credit)$
13:    $b\leftarrow b+credit$; $sb_{i}\leftarrow sb_{i}+credit$; $rem_{i}\leftarrow rem_{i}-credit$
14:end procedure
4.3 Congestion Control Algorithm
Algorithm 1 describes the behavior of a SIRD receiver. When credit is available in the global bucket and on the command of the credit pacer (§4.4), the receiver tries to allocate credit to an active sender (ln. 8). It first selects one of the senders with available credit in the per-sender bucket (ln.9) based on policy (ln.10). The receiver sends a CREDIT packet (ln. 12) and reduces the available credit in the global and per-sender buckets, and updates the remaining required credit for that message (ln. 13). Whenever a DATA packet arrives (ln. 1), if it is scheduled, the receiver replenishes credit in the global bucket (ln. 3) and the per-sender bucket (ln. 4). Then, it executes the two independent AIMD control loops to adapt to a congested sender (ln. 5) and a congested core network (ln. 6), respectively. The receiver sets the size of the per-sender bucket as the minimum of the two values (ln. 9).